OpenStudy (rainbow_rocks03):

On a map 1 inch = 5 miles. If a road is 130 miles long, how long will the road on the map be? ______ inches

OpenStudy (rainbow_rocks03):

Will medal+fan for the best response , no direct answers I want the explanation.

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }=\frac{ x }{ 130 }\] Just cross multiply and then get x alone.

OpenStudy (rainbow_rocks03):

I think the answer is 650 because

OpenStudy (rainbow_rocks03):

130*5=650

OpenStudy (anonymous):

That would be wrong. You are treating it as if every mile is 5 inches. It is the opposite.

OpenStudy (rainbow_rocks03):

oh so 130/5 ?

OpenStudy (rainbow_rocks03):

so is my answer 26 ?

OpenStudy (anonymous):

Yes

OpenStudy (rainbow_rocks03):

thx for the help. :)

OpenStudy (anonymous):

No problem :)